OpenStudy (anonymous):

If 20 waves of water pass you in 5.0 s and the waves are 2.0 m long, how fast are they moving? 1.0 m/s , 2.0 m/s , 4.0 m/s , or 5.0 m/s ?

OpenStudy (michele_laino):

is 2.0 meters the wave length ?

OpenStudy (anonymous):

i think so :)

OpenStudy (anonymous):

ohh so our solution is 2.0 m/s?

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

if 20 waves pass in 5 seconds, the one wave pass in 5/20=0.25 seconds

OpenStudy (michele_laino):

so I see one wave length every 0.25 seconds, and the requested speed is: 2/0.25=8

OpenStudy (michele_laino):

what is wrong in my reasoning?

OpenStudy (anonymous):

ohh okay! ermm i am not sure... :/ the solution is 8 though right? :/

OpenStudy (anonymous):

i am not sure why that is not a choice? :/

OpenStudy (anonymous):

because your reasoning makes sense :/

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay! thank you!! :)

OpenStudy (michele_laino):

I'm sorry!

OpenStudy (anonymous):

haha it is okay! thank you:)

OpenStudy (michele_laino):

please try to ask to another tutor

OpenStudy (anonymous):

okie:)

OpenStudy (anonymous):

thankyou!!

OpenStudy (michele_laino):

please wait I call for another tutor

OpenStudy (anonymous):

okie! :)

OpenStudy (michele_laino):

@Australopithecus please help!

OpenStudy (michele_laino):

@Hero please help

OpenStudy (michele_laino):

@amistre64 please help

OpenStudy (michele_laino):

@sammixboo please help

OpenStudy (michele_laino):

@Jaynator495 please help

OpenStudy (michele_laino):

@Nnesha @ParthKohli please help

Parth (parthkohli):

\[v = f \times \lambda\]

Parth (parthkohli):

v - velocity of the wave. f - frequency of the wave (number of waves passing in 1 second). \(\lambda\) - wavelength, or length of a wave.

OpenStudy (anonymous):

so f=20/5 = 4 waves per second? and wavelength is 2.0 m ?

Parth (parthkohli):

That is correct. The answer is thus \(4 ~s^{-1} \times 2 ~m = 8 ~ms^{-1}\).

OpenStudy (anonymous):

ooh okay! hmm i am not sure then :/ i was not given that choice :( i think maybe a typo? haha :P but yes, thank you both very much!!:)

OpenStudy (michele_laino):

thanks! @ParthKohli